OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x))

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)).

OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4

OpenStudy (cggurumanjunath):

http://openstudy.com/study#/updates/527e631ae4b022369a7c1ed1

OpenStudy (amistre64):

f, composed of g f(g) = 9g-2 but since g = -x+3, replace g f(g(x)) = 9(-x+3)-2

OpenStudy (anonymous):

r u sure its not -9x - 2

OpenStudy (amistre64):

pretty sure .... im fairly confident that it is exactly like a posted :/

OpenStudy (lukecrayonz):

Think of it as f with g inside of it. f(g(x)) So everytime you see x inside of f(x), input g instead. 9x-2 input g(x) for x 9(-x+3)-2 distribute 9 -9x+25

OpenStudy (anonymous):

Yes, it is exactly as @amistre64 explained it.

OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (lukecrayonz):

Same exact thing. Literally.

OpenStudy (anonymous):

what about Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (lukecrayonz):

Okay seriously just learn it. f(g(x)) means put g(x) inside of f(x) g(f(x)) means put f(x) inside of g(x) So when you see x, put the coordinated term inside of it. When you see f(1), it means put 1 for everytime you see x So now, it's saying f(g(x)) so put g(x) everytime you see x. We could write it as f(10x-6) but it just looks messy, so we write it as f(g(x)). So, take f(x) = -4x + 7 and put g(x) inside of it everytime you see x.

OpenStudy (lukecrayonz):

-4(10x - 6)+7 -40x+24+7 24+7=31 -40x+31

OpenStudy (anonymous):

tyvm